### v1

# DeVAIC: A Tool for Security Assessment of AI-generated Code

Abstract
Context: AI code generators are revolutionizing code writing and software development, but their training on large datasets, including potentially untrusted source code, raises security concerns. Furthermore, these generators can produce incomplete code snippets that are challenging to evaluate using current solutions.
Objective: This research work introduces DeVAIC (Detection of Vulnerabilities in AI-generated Code), a tool to evaluate the security of AI-generated Python code, which overcomes the challenge of examining incomplete code.
Method: We followed a methodological approach that involved gathering vulnerable samples, extracting implementation patterns, and creating regular expressions to develop the proposed tool. The implementation of DeVAIC includes a set of detection rules based on regular expressions that cover $35$ Common Weakness Enumerations (CWEs) falling under the OWASP Top $10$ vulnerability categories.
Results: We utilized four popular AI models to generate Python code, which we then used as a foundation to evaluate the effectiveness of our tool. DeVAIC demonstrated a statistically significant difference in its ability to detect security vulnerabilities compared to the state-of-the-art solutions, showing an $F_{1}$ Score and Accuracy of $94\%$ while maintaining a low computational cost of $0.14$ seconds per code snippet, on average.
Conclusions: The proposed tool provides a lightweight and efficient solution for vulnerability detection even on incomplete code.
keywords:
Static code analysis, Vulnerability detection, AI-code generators, Python
[…]
organization=University of Naples Federico II, city=Naples, postcode=80125, country=Italy
1 Introduction
We live in an era where AI-code generators are revolutionizing the process of writing code and software development. AI-powered solutions like GitHub Copilot copilot , OpenAI ChatGPT chatGPT , Google Gemini bard , and Microsoft Copilot bing have shown their ability to translate into programming code what users request with natural language (NL) descriptions (e.g., English language).
The effectiveness with which AI-code generators produce code has brought users of different levels of skills and expertise to adopt such solutions to promptly solve programming problems or to integrate AI-generated code into software systems and applications. The other side of the coin is that their widespread usage is out of any quality control, leading to a question to preserve the security of the software development process: can we trust the AI-generated code? Ergo, from a software security perspective, is the code generated by AI secure and free of software vulnerabilities?
The concern stems from the consideration that these solutions are trained on a large amount of publicly available data. For instance, GitHub Copilot utilizes training data that consists of an extensive amount of source code, with billions of lines collected from publicly accessible sources, including code from GitHub’s public repositories github_blog . Unfortunately, as often happens, quantity does not coexist with quality. In fact, the huge amount of data used to train the AI-code generators may include deprecated functions and libraries, which may lead to the exploitation of vulnerabilities when adopted, or it could intentionally have buggy or insecure code used to poison the code generators during the training phase improta2023poisoning ; li2023poison ; hussain2023trojanedcm ; hamer2024just .
This issue is not unexpected at all. It suffices to think that users are always warned to exercise caution when using the AI-generated code (e.g., “The users of Copilot are responsible for ensuring the security and quality of their code” warn_copilot , “Use the code with caution” bard_blog , “This code is generated by artificial intelligence. Review and use carefully” bing_blog ).
However, it is not clear what users can do to assess the security of the AI-generated code to integrate it into their code base. Manual analysis, the go-to method for security experts, becomes unfeasible due to the volume and rate deployment of AI-generated code liguori2021evil ; liguori2022can . In fact, the speed at which these solutions operate can overwhelm even the most experienced security professionals, making it difficult to thoroughly review each line of code for potential vulnerabilities.
To perform code vulnerability detection, state-of-the-art solutions provide several static analysis tools. Such tools usually parse the Abstract Syntax Tree (AST) of the code ma2022code ; bandit ; pyt ; codeql , a hierarchical representation of its structure, and require complete programs to check whether the code contains software vulnerabilities. However, since AI-code generators are fine-tuned on corpora which often contain samples of code, i.e., code snippets humaneval ; django ; liguori-etal-2021-shellcode ; conala , they often do not produce complete programs liguori2022can ; zhong2023study , making infeasible the application of these tools in this context.
A different solution is represented by tools implementing pattern-matching approaches to detect software vulnerabilities mastjik2015comparison ; patel2015implementation ; walenstein2007exploiting . These tools require users to set up ad-hoc configuration files to specify the vulnerabilities they wish to detect via matching patterns mordahl2023automatic ; nadi2014mining , hereby requiring a non-trivial manual effort and limiting their adoption in practice.
Previous studies overcome the limitations of static analysis tools by adopting AI-based solutions serving as vulnerability detectors. These fully automated solutions do not require any human effort and can analyze incomplete programs, but they may be susceptible to a high rate of false alarms chakraborty2021deep ; chen2023diversevul ; li2021vulnerability ; li2018vuldeepecker . The existing approaches using AI for vulnerability prediction face challenges related to training data and model choices. Furthermore, it has been shown that AI models struggle to understand the complexity of code structures and identify security issues fang2023large ; al2023chatgpt ; cheshkov2023evaluation .
All the above limitations make evident the existence of a gap in the automatic security assessment of the AI-generated code. Addressing these limitations is crucial for enhancing the trustworthiness of AI-generated code and ensuring the security of the systems built upon it.
This paper presents DeVAIC (Detection of Vulnerabilities in AI-generated Code), a tool that performs static analysis of Python code by implementing a set of detection rules. More precisely, the tool uses a set of regular expressions that cover $35$ Common Weakness Enumeration (CWE), a community-developed list of software and hardware weakness types, to detect vulnerabilities in Python code, one of the most used programming languages octoverse ; meyerovich2013empirical ; statista . The tool does not require the completeness of the code, making it suitable to detect and classify vulnerabilities in AI-generated code snippets, nor to create ad-hoc detection rules, hence overcoming the limitations of state-of-the-art solutions.
We used DeVAIC to detect vulnerabilities in the code generated by four well-known public AI-code generators starting from NL prompts. Our experiments show that the tool automatically identifies vulnerabilities with $F_{1}$ Score and Accuracy both at $94\%$ and low computational times ($0.14$ seconds for code snippet, on average). Also, we show that DeVAIC exhibits performance superior to the state-of-the-art solutions, i.e. CodeQL codeql , Bandit bandit , PyT pyt and Semgrep semgrep , that are widely used to perform vulnerability detection of the code pearce2022asleep ; dunlap2023finding ; bakhshandeh2023using ; kathikar2023assessing .
In the following, Section 2 discusses related work; Section 3 introduces a motivating example; Section 4 presents DeVAIC; Section 5 describes the results; Section 6 discusses the threats to validity; Section 7 concludes the paper.
2 Related Work
Static analysis is one of the most commonly used techniques to detect vulnerabilities in the code chess2004static ; li2017static ; pan2020systematic ; goseva2015capability .
The state of the art provides several static analysis tools for checking security issues in the code, e.g., Python-specific tools like Bandit bandit and PyT pyt , or multilanguage ones such as Semgrep semgrep and CodeQL codeql , which are widely used to detect vulnerabilities within the code pearce2022asleep ; dunlap2023finding ; bakhshandeh2023using ; kathikar2023assessing ; ruohonen2021large ; kapustin2023static ; peng2019python ; lyons2021meta ; bandara2020fix ; gobbi2023poster ; cherry2022static .
Except for Semgrep, these tools require working on complete code due to the preliminary modeling of AST from the code under examination. The vulnerability detection is made by running appropriate plugins (for Bandit and PyT) or queries (for CodeQL) against the AST nodes. For code that consists of snippets rather than complete programs, these analysis tools cannot define the AST, thus having any possibility for conducting their detection analyses. Semgrep semgrep is a static analysis tool that uses a pattern-matching approach and that does not require the AST modeling of code before running the detection rules. To detect vulnerabilities, users need to configure and customize the tool by writing regex patterns in a configuration file. However, the limitation is that we cannot assume in advance that all users can write accurate regex patterns. Moreover, this approach could polarize vulnerability hunting, focusing on those the user believes to find, potentially overlooking others that are effectively present mordahl2023automatic ; nadi2014mining . For this reason, as often happens, this type of solution offers a set of rules publicly available for the scanning of the code under examination semgrep_registry .
To overcome the limitations of static analysis tools, previous work investigated the use of AI to perform vulnerability detection, using it as a static analyzer. One of the benefits of their adoption is that they can analyze incomplete code. However, previous work shows that the outcomes of their detection exhibit numerous false assessments.
For instance, Chen et al. chen2023diversevul released a new vulnerable source code dataset to assess state-of-the-art deep learning methods in detecting vulnerabilities. The authors show the limited performance of these solutions in terms of high false positive rates, low $F_{1}$ Scores, and difficulty in detecting hard CWEs. Fang et al. fang2023large crafted a dataset that pairs real-world code with obfuscated versions, which they used as input for large language models (LLMs) to evaluate their ability to analyze input code samples and test if LLMs can be employed for defensive analysis tasks. The study found that larger models are able to comprehend and explain unobfuscated code, whereas smaller models failed in this task. However, the model’s understanding of the code was limited when working with obfuscated code. Al-Hawawreh et al. al2023chatgpt and Cheshkov et al. cheshkov2023evaluation evaluated the performance of the ChatGPT for detecting vulnerabilities in code, finding that ChatGPT’s results still needed careful monitoring. Khoury et al. khoury2023secure explored how safe is the code generated by ChatGPT, using this model for both code generation and assessment. Their analysis outlined that, despite ChatGPT’s awareness of potential vulnerabilities, it still can produce unsafe code. Instead of using AI-based solutions, Sandoval et al. sandoval2023lost  sandoval2022security assessed the security of AI-generated code by performing a manual code analysis. Although manual code checking has advantages, its susceptibility to human error, scalability problems and subjectivity emphasises the importance of using automated tools for a more consistent code analysis.
This work presents a tool to perform vulnerability detection of Python code. The tool does not require complete programs to analyze as it also works for code snippets, making it suitable for assessing the security of AI-generated code. Similar to Semgrep, the proposed tool implements a pattern-matching approach while employing a lesser quantity of rules and achieving better results in terms of Recall, $F_{1}$ Score, and Accuracy. Indeed, it does not require any human effort to configure detection rules, hence overcoming the limitations of previous static analysis tools.
The tool can identify vulnerabilities belonging to the OWASP categories with the discovery of $35$ CWEs, a coverage level comparable with other research work. For instance, Hajipour et al. hajipour2023systematically consider $13$ CWEs for the assessment of the security risks posed by code language models through an approximation of the inversion of the black-box models using few-shot prompting; Pearce et al. pearce2022asleep , moreover, refer to the MITRE’s Top $25$ CWE list to assess the code generated by Copilot. CodeQL provides queries to assess the security of Python code with a coverage of $44$ CWEs codeql_coverage . Finally, the tool is able to detect vulnerabilities in the AI-generated Python code with a low rate of false positives and negatives, overcoming the performance of other state-of-the-art tools used as static code analyzers, and low computational cost.
3 Motivating Example
AI models for code generation have shown significant advancements, largely attributed to their intuitive, user-friendly interfaces. These models enable users to request code generation through NL descriptions, specifying both the programming language and the level of detail needed for the task at hand. For illustrative purposes, consider the following example:
Listing 1: Example of NL prompt.
Write a Python code that defines a YAML payload from a request,
loads this payload into a dictionary called prod_and_qt
by using a function for loading.
Typically, AI models can produce either complete or partial snippets of code based on the nature of the training data they have been exposed to. An example output for the aforementioned prompt might look like this:
Listing 2: Example of code generated by AI models.
def load_yaml(yaml_payload):
   prod_and_qt = yaml.load(yaml_payload)
   return prod_and_qt
yaml_payload = ””
product:
name:iPhone
quantity:10
prod_and_qt = load_yaml(yaml_payload)
print(prod_and_qt)
The code generated by the model is what is requested in the NL prompt. However, the code is vulnerable due to the yaml.load() function used to process YAML data from yaml_payload, which may contain untrusted content. Indeed, if manipulated by a malicious user, this payload may include dangerous data or code. The vulnerability is in the execution of the yaml.load() function, because it interprets the YAML payload as Python code, enabling the possible execution of malicious instructions if the code lacks a proper validation, like in this case. The CWE associated with this vulnerability is CWE-502, commonly known as Deserialization of Untrusted Data, and related to the Software and Data Integrity Failures category of OWASP’s Top 10. A simple way to address this issue is reported by the official PyYAML documentation, which recommends using yaml.safe_load() to read YAML from unreliable and untrusted sources PyYAML_doc .
In most cases, users are unaware of software vulnerabilities and may not be able to manually verify the security of the code, thereby including the produced outcome into an existing codebase. This issue is further exacerbated by the fact that state-of-the-art static code analyzers, such as CodeQL, Bandit, PyT, etc., do not generate the report for this specific code snippet. In fact, the code reported above is incomplete due to the lack of the import statement at the beginning of it, making these tools unable to perform the vulnerability detection, as explained in Section 2. This code characteristic does not consent to the modeling of the AST as it lacks the necessary dependency (i.e., import yaml) for the invoked API (i.e., the yaml.load() function). Conversely, Semgrep’s textual analysis approach did examine the code but resulted in a False Negative (FN), highlighting the challenge of ensuring the security of AI-generated code.
These issues underscore the imperative to critically evaluate AI-generated code for security vulnerabilities, thereby ensuring the safe integration of such code into larger systems.
4 DeVAIC Workflow
To overcome the issues described in Section 3, we present DeVAIC, a tool to detect vulnerabilities in Python code. The tool does not require the completeness of the code, making it suitable for AI-generated code.
The DeVAIC tool works as a text scanner and does not require the AST modeling for the code analysis, employing regular expressions to identify vulnerable code patterns. However, crafting effective regex requires a deep understanding of the implementation patterns we want to find.
Figure 1: The DeVAIC workflow.
To this aim, we extensively reviewed the literature to collect datasets of unsafe Python code with the associated information of the implemented CWE [^1] The Common Weakness Enumeration (CWE) is a list of common software and hardware weakness types published annually by the Massachusetts Institute of Technology Research and Engineering® (MITRE) Corporation., extracting only the snippets that implement the most recent and critical weaknesses identified in OWASP’s Top $10$ Vulnerabilities [^2] The Open Web Application Security Project® (OWASP) draws up the Top $10$ Application Security List every four years to define the $10$ most prevalent web application vulnerabilities. report. Since OWASP categories encompass various CWEs that share the same vulnerability typology, we employ these categories to cluster snippets together for further detailed analysis (see § 4.1).
Using a named entity tagger, we standardize snippets by replacing variable names, and input and output parameters of functions to reduce the variability of the code. Then, we compare snippet pairs within each OWASP category based on their similarity level, facilitating automated identification of common implementation patterns using the Longest Common Subsequence (LCS), i.e., the longest sequence that appears as a (not necessarily contiguous) subsequence in both snippets. By using the identified patterns, we infer regex-based detection rules able to identify vulnerabilities with similar patterns and assess their corresponding vulnerability types (see § 4.2).
Finally, we collect all the detection rules in a bash script that takes an input file containing line-by-line code to analyze. After scanning the file, the tool generates a report of the detection, including the number of snippets identified as vulnerable and their category according to the OWASP (see § 4.3).
Figure 1 shows a detailed flowchart outlining the steps through which we developed the tool. In the rest of this Section, we detail the steps of the workflow.
4.1 Vulnerable Code Collection
In our pursuit to collect vulnerable snippets, we performed a comprehensive study of the literature. State-of-the-art provides numerous corpora containing vulnerable code, some aimed at poisoning models to induce them to generate vulnerabilities, while others serve to evaluate AI models as vulnerability detectors cotroneo2023vulnerabilities ; chen2023diversevul .
Since we target Python code, we selected two corpora suitable to our purpose, discarding those containing snippets written in different programming languages (e.g., DiverseVuln chen2023diversevul , Big-Vul bigvul ; fan2020ac , SARD sard ):
  1. 1.
    SecurityEval seceval : It is a dataset built to evaluate the security of code generation models and to compare the performance of some state-of-the-art static analysis tools, such as Bandit and CodeQL siddiq2022seceval . The dataset contains $130$ Python code samples covering $75$ vulnerability types, which are mapped to prompts written in Natural Language (NL).
  2. 2.
    Copilot CWE Scenarios Dataset datasetCWE : It encompasses the source code utilized to craft NL prompts for generating code with Large Language Models (LLMs) llmseceval ; llmseceval2023 . It covers code associated with $89$ distinct vulnerable scenarios.
Table 1: List of $35$ selected CWEs. In blue we indicate the CWEs belonging to at least one of the MITRE’s top $25$ of the last three years. We use “-” to indicate the absence of a specific CWE in the relative top $40$.
OWASP 2021 CWE Rank MITRE 2021 Rank MITRE 2022 Rank MITRE 2023
Broken Access Control CWE-022
8
8
8
CWE-377
CWE-425
CWE-601
37
35
32
Cryptographic Failures
CWE-319
35
39
CWE-321
CWE-326
CWE-327
CWE-329
CWE-330
CWE-347
CWE-759
CWE-760
Identification and Authentication Failures
CWE-295
26
26
34
CWE-384
Injection CWE-020
4
4
6
CWE-078
5
6
5
CWE-079
2
2
2
CWE-080
CWE-090
CWE-094
28
25
23
CWE-095
CWE-096
CWE-099
CWE-113
CWE-116
CWE-643
CWE-1236
Insecure Design
CWE-209
CWE-269
29
29
22
CWE-434
10
10
10
Security Logging and Monitoring Failures (SLMF)
CWE-117
Security Misconfiguration CWE-611
23
24
28
Server-Side Request Forgery (SSRF)
CWE-918
24
21
19
Software and Data Integrity Failures (SDIF)
CWE-502
13
12
15
To develop detection rules that address the most recent and critical weaknesses, we extracted from these corpora only the vulnerable code snippets that implement the CWEs present in the most recent OWASP Top $10$ Vulnerabilities report from 2021 owasp_2021 ; honkaranta2021towards . Overall, we selected $240$ vulnerable snippets that implement $35$ CWEs related to $9$ OWASP categories [^3] MITRE itself is the authority responsible for establishing the correlation between CWEs and the OWASP Top $10$. Among the selected CWEs, $10$ of them can be found in at least one of the MITRE’s Top $25$ lists from the past three years top40_2021 ; top40_2022 ; top25_2023 . The use of $35$ CWEs is consistent with other detection setups used in previous research studies hajipour2023systematically ; pearce2022asleep . Table 1 shows the selected CWEs with their OWASP category.
4.2 Detection Rules Implementation
To define detection rules, we search for common patterns among the selected code snippets based on the assumption that similar vulnerabilities manifest in similar ways thomas2020learning ; honkaranta2021towards ; mastjik2015comparison . Pattern-matching approaches such as LCS, N-gram and others, have been widely used to identify the same family malware mastjik2015comparison ; walenstein2007exploiting ; patel2015implementation ; zhang2022example . In our work, we applied the LCS algorithm to identify vulnerable patterns among snippets belonging to the same OWASP category.
As an example, consider the pair of code snippets shown in Table 2 - “Original Snippet” column, which belongs to the Injection OWASP category (i.e., user-supplied data is not validated, filtered, or sanitized by the application). Specifically, the code snippet in Row #1 addresses CWE-020 (Improper Input Validation), whereas the snippet in Row #2 is linked to CWE-080 (Basic XSS). Despite implementing distinct CWEs, these samples share analogous implementation patterns, i.e., the use of an output value from a function of a Python module without any sanitization, subsequently used as an input parameter in a different function.
Table 2: Examples of standardization for code snippets. The common longest subsequences of original snippets and the standardized snippets are in blue.
Original Snippet Standardized Snippet
S1
def temperature_for_location(): file_yaml = request.args.get("file", None) json_payload = yaml.load(file_yaml) def temperature_for_location(): var0 = request.args.get(var1, var2) var3 = yaml.load(var0)
S2
name = request.args.get(’name, ’’) response = make_response(name) return response
var0 = request.args.get(var1, var2) var3 = make_response(var0) return var3
Therefore, to search for similar code snippets, we first check the similarity for each pair of snippets grouped in the same OWASP category. To this aim, we use SequenceMatcher, a class of the Python module difflib, that computes the Longest Common Subsequence (LCS) with no junk elements, i.e., tokens that are not considered in the matching (e.g., the newline character \$\backslash n$), among different code snippets. First, we compute the similarity ratio between two code snippets, ranging from 00 (total mismatching) to $1$ (perfect matching). Based on empirical observations, we focus exclusively on pairs exhibiting a similarity ratio exceeding $50\%$ rao2018efficient ; rao2018characteristic ; rao2018partial . This decision is grounded in the assumption that when two snippets share at least half of their content, we can infer the existence of a meaningful common pattern. Then, we find common patterns among snippets that meet the similarity threshold within each OWASP category by computing the LCS ma2017vurle ; mastjik2015comparison .
To support the LCS in finding common patterns, we standardize all the snippets for each OWASP category, i.e., we reduce the randomness of the code snippets, by using a named entity tagger, which returns a dictionary of standardizable tokens for the input and output parameters of functions, extracted through regular expressions. We replace the selected tokens in every intent with “var#”, where # denotes a number from 00 to $|l|$, and $|l|$ is the number of tokens to standardize. This data processing method prevents snippets that have a low similarity score due to parameters (e.g., parameter names containing a high number of tokens) from not being grouped to find common patterns.
Table 2 - “Standardized Snippet” column shows the standardization process performed on the snippets, where the parameters of the original code snippets are replaced (e.g., the input parameter “file” is replaced with var1 in the code snippet of row #$\#1$). Due to the standardization, the similarity ratio increases from $\sim 37\%$ (obtained on the original snippets) to $\sim 63\%$. As a result, the LCS returns a clearer and easier-to-identify pattern, as shown in the table.
To identify patterns, we create detection rules by using regular expressions (regex) that, by their nature, operate on patterns within the text (code). As they do not require the completeness of the code to identify specific patterns, they are well-suited for analyzing incomplete or partial programs, which is often the case of AI-generated code. For example, a detection rule based on LCS result of the example in Table 2 can involve detecting the input function request.args.get(), whose output (i.e., var0) lacks proper sanitization and is subsequently utilized as an input parameter for another sink function (i.e., passing the variable var0 as a parameter in brackets, regardless of the specific function in which it is passed). Since rules are created by patterns found in the same OWASP category, each rule is associated with a category. Therefore, when the tool analyzes a code snippet containing the previous pattern, the rule detects the vulnerability and indicates the related OWASP category.
In the implementation of detection rules, we adopt domain-specific language designed for text processing such as awk, grep, and sed, which are standard features of most Unix-like operating systems. For example, to implement a rule to detect the pattern of the previous example, we use awk to store the name of the variable that contains the output of the request.args.get() function. Then, we employ grep with a specific regular expression that checks if this variable is used in subsequent code without any sanitization.
Overall, we implement $85$ detection rules that cover the $35$ CWEs over the $9$ OWASP categories shown in Table 1. Since there is no unique way to implement a CWE, the number of detection rules overcomes the number of CWEs.
4.3 Tool Execution
As DeVAIC uses standard features of most Unix-like operating systems, it is highly portable. It takes as input a file in TXT format containing a set of code snippets, each written on a single line. Multi-line code snippets (e.g., a function) are separated by the newline character \$\backslash n$.
While scanning the entire set of snippets in the input file, DeVAIC executes all the detection rules to detect any vulnerabilities. If the tool identifies a vulnerability through a rule, the tool continues the execution of all the detection rules since the same snippet may implement different vulnerabilities, even belonging to different OWASP categories. At the end of the execution, the tool returns in output a file containing a summary report of the detection results. The report includes:
  • 1.
    The number of code snippets analyzed by the tool;
  • 2.
    The number/percentage of code snippets identified as safe, i.e., that do not contain any vulnerability, and the number/percentage of code snippets identified as unsafe along with the classification of the vulnerabilities according to OWASP Top $10$;
  • 3.
    The overall execution time on the entire dataset of snippets and the average execution time per single snippet.
The tool also produces a file that exhibits the detection results, highlighting the OWASP categories for each vulnerable snippet. Since a single code may encompass multiple vulnerabilities from different categories, the occurrences of snippets that fall in the various OWASP categories may exceed the total number of unsafe code snippets.
5 Experimental Evaluation
We evaluate the effectiveness of our tool, DeVAIC, through experiments conducted on code generated by four distinct publicly available AI models, i.e., Google Gemini (LaMDA’s successor), Microsoft Copilot (GPT-4), OpenAI ChatGPT (GPT-3.5), and GitHub Copilot (GPT-4). These models are accessible via APIs and generate code suggestions starting from NL prompts.
5.1 NL Prompts Details
We used a set of $125$ NL prompts for each of the mentioned models to generate Python code, for a total of $500$ code snippets used to assess DeVAIC’s detection skills.
We designed these prompts starting from those provided by three publicly available datasets about code generation, i.e. SecurityEval SecurityEval_prompts , LLMSecEval LLMSecEval_prompts and CodeXGLUE CodeXGLUE_prompts . First of all, we modified these prompts by avoiding the insertion of implementation details. This line of action allows the model to generate code according to their discretion without any influence on how to implement a task. Moreover, we adopted the perspective of an ordinary, nonmalicious user who requested the model to generate code for a given task, which means that our prompts do not contain explicit triggers to induce the prediction of unsafe code.
The number of tokens for these $125$ prompts has an average value of $15.5$ (median value is $14$). Typically, most prompts fall within the range of $11$ to $19$ tokens, with a few outliers presenting a minimum of $7$ and a maximum of $34$ tokens. The variation in prompt length is attributed to the need for additional details to clarify the request, although it is important to remark that prompts do not contain any implementation details.
Table 3: Examples of vulnerable code produced by models. The vulnerable part of the code is highlighted in red.
NL prompt Unsafe prediction
Write a Python code with a get_name() function that gets the name parameter from the URL by using request module by importing flask and opens the file names/name returning a file object
[…]
Get the user input from the request and eval it
[…]
Table 3 displays two examples of NL prompts and a sample of vulnerable code generated by the models we employed. Row #1 showcases a model’s prediction that uses the request.args.get() function of the Flask module. The only validation produced in the code (i.e., the only one defined by the user in the prompt) is to verify if the output parameter of the request.args.get() function is an empty variable. Nevertheless, it is not sufficient because, even if this parameter is not empty, it may contain dangerous characters (i.e., “../”, “../..”, etc.) that can allow a malicious user to access files or directories outside of the intended directory (i.e., name/). The CWE associated with this kind of vulnerability is CWE-022, commonly known as “Path Traversal”. This CWE is related to the Broken Access Control category of OWASP’s Top 10. To enhance the security, the code has to ensure that the file path is actually within the expected directory before opening the file itself owasp_pathtrav . We also notice that the generated code snippet lacks completeness, making the application of some static analysis tools not feasible.
Finally, Row #2 shows that the request to evaluate a user input involved the generation of code containing the eval() function. This function is considered dangerous due to the potential injection of malicious code and is related to CWE-095, also known as “Eval Injection”, which belongs to the Injection OWASP category. In fact, the official ast documentation recommends using literal_eval(), a more secure alternative to eval(), that can evaluate only a restricted set of expressions ast_doc .
5.2 Manual Analysis
After submitting the NL prompts, the models generated a total of $500$ code snippets. The average number of tokens for the code is $54$ (with a median of $42$, a maximum of $205$, and a minimum of $4$), with most of the code snippets ($254$ in total, $\sim 51\%$) falling within the range of $22$ to $88$ tokens. Furthermore, as shown in Figure 2, every model produced some instances of incomplete code, i.e., the model provided code functions without the necessary import statement at the beginning. For each group of $125$ snippets produced by each model, we reported $8$ incomplete code snippets for Google Gemini ($\sim 6\%$), $12$ for Microsoft Copilot ($\sim 10\%$), $39$ for GitHub Copilot ($\sim 31\%$) and $6$ for OpenAI ChatGPT ($\sim 5\%$), reaching a total of $13\%$ of incomplete code on $500$ code generated. This result underlines the importance of evaluating AI-generated code with a tool able to overcome this issue.
Figure 2: Occurrences of complete and incomplete code generated by each model.
After grouping the code for each model and converting them in a TXT file with snippets written line by line, we ran DeVAIC to check for vulnerable implementation patterns in the outcomes. The assessment of the detection results needs manual inspection, involving the evaluation of each code procured by models in terms of True Positives (TPs), False Positives (FPs), True Negatives (TNs), and False Negatives (FNs). More precisely, we have a TP case when both DeVAIC and manual analysis detect a software vulnerability; similarly, when both DeVAIC and manual analysis do not detect any vulnerability in the code, then we have a TN case. When DeVAIC identifies a vulnerability not confirmed by manual inspection, then it is an FP case. Conversely, when the tool does not identify a vulnerability within the code but manual analysis does, then it is an FN case.
As manual classification can be susceptible to errors, three researchers of our research group, who have skills in Python code vulnerability detection, conducted this evaluation independently. Subsequently, they compared their evaluation and performed an in-depth analysis of the few discrepancy cases. This collaborative effort resulted in a complete alignment, achieving a $100\%$ consensus in the final evaluation.
On average, the four models produced $54\%$ of vulnerable code ($271$ vulnerable code over $500$ predictions), which allows us to comprehensively assess the tool’s detection ability in both vulnerable and non-vulnerable instances, providing a robust and fair measure of their overall performance across various scenarios.
5.3 Experimental Results
We assessed the tool’s detection ability by using standard metrics typically employed in this field, i.e., Precision, Recall, $F_{1}$ Score, and Accuracy. We computed these metrics with the TPs, TNs, FPs, and FNs identified during the manual analysis (see § 5.2).
Table 4: Evaluation of detection results comparing DeVAIC with the state of the art.
Detection Tool Google Gemini OpenAI ChatGPT-3.5 Microsoft Copilot GitHub Copilot All models
Precision
DeVAIC
0.97
1.00
0.95
0.95
0.97
Bandit
0.89
0.81
0.83
0.82
0.84
CodeQL
0.79
0.86
0.83
0.95
0.85
Semgrep
0.87
0.98
0.94
0.85
0.91
PyT
1.00
0.89
1.00
1.00
0.96
Recall
DeVAIC
0.95
0.96
0.90
0.86
0.92
Bandit
0.70
0.69
0.51
0.58
0.62
CodeQL
0.36
0.54
0.42
0.26
0.39
Semgrep
0.55
0.69
0.58
0.51
0.58
PyT
0.11
0.11
0.08
0.04
0.09
$F_{1}$ Score
DeVAIC
0.96
0.98
0.92
0.90
0.94
Bandit
0.78
0.74
0.63
0.68
0.72
CodeQL
0.49
0.67
0.56
0.41
0.54
Semgrep
0.67
0.81
0.72
0.64
0.71
PyT
0.20
0.20
0.16
0.08
0.16
Accuracy
DeVAIC
0.95
0.98
0.93
0.89
0.94
Bandit
0.77
0.74
0.71
0.69
0.73
CodeQL
0.56
0.70
0.68
0.58
0.63
Semgrep
0.69
0.82
0.78
0.67
0.74
PyT
0.48
0.50
0.56
0.46
0.50
To provide context for the evaluation, we compared DeVAIC’s performance with a baseline. In our analysis, we utilized CodeQL version v2.16.4 and the two Security test suites sec_testsuite2 ; sec_testsuite1 of queries for Python. We also used Bandit version 1.7.7, Semgrep version 1.61.1, and python-taint module version 0.42, also known as PyT. We remark again that several snippets generated by the models were incomplete, preventing us from directly analyzing them with some well-known static analysis tools, such as Bandit, CodeQL, or others. At first, we analyzed the model predictions with DeVAIC. Then, we modified the incomplete ones ($65$ in total) by inserting the import statement when lacking to compare DeVAIC’s evaluation results with the state of the art.
Considering the average values for all models, Table 4 shows that DeVAIC achieved high values for each metric. Regarding Precision, DeVAIC reaches $97\%$, which is comparable with the values achieved by Semgrep [^4] When we used Semgrep, we executed a total of 1291 rules from the official ruleset registry for Python semgrep_registry . ($91\%$) and PyT ($96\%$). Furthermore, DeVAIC achieved a maximum Recall of $92\%$ on average, surpassing the baseline with a considerable gap compared to the employed solutions, which obtain a maximum of $62\%$ (i.e., Bandit). Finally, DeVAIC shows superior performance to other state-of-the-art solutions ($\geq 20\%$) with average values of $94\%$ for both $F_{1}$ Score and Accuracy.
We further evaluated if, for each metric (i.e., Precision, Recall, F1 Score, and Accuracy), there is a statistical difference between DeVAIC and the state-of-the-art solutions using the non-parametric Wilcoxon rank sum test. The null hypothesis is that the two samples derive from the same population (i.e., the two populations have equal medians). If the null hypothesis is rejected, the Wilcoxon rank sum test indicates that the two samples are statistically different. As deducible by the considerations explained above, for the Precision, DeVAIC is statistically different only with Bandit while obtaining optimum results for all of the other metrics, for which DeVAIC results are statistically different (and better) with every other result obtained with the state of the art. The previously mentioned considerations highlight the outstanding performance of DeVAIC.
Figure 3: Analysis of snippet lengths variability for each classification. The median values are in orange, while the mean values are in green.
Then, we performed an in-depth analysis of the results provided by our tool to investigate the cases of FN and FP. Since DeVAIC behaves like a text scanner, we first checked whether the performance of the tool is affected by the complexity, in terms of number of tokens, of the code to analyze. Figure 3 shows four boxplots to illustrate the complexity of the code across the TP, TN, FP, and FN cases. The figure highlights that the interquartile ranges, i.e., the height of the boxplots, are greater for TP and TN cases ($59$ and $57$, respectively) than the FN ($46$) and FP cases ($29$). Moreover, the average number of tokens per category, which is $46$ for TP, $40$ for TN, $44$ for FP, and $43$ for FN, proves that the cases of misclassification do not depend on the complexity of snippets as DeVAIC is effective in detecting vulnerabilities even for complex code snippets.
5.4 Computational Cost
We analyzed the computational times of DeVAIC. We run the tool on a computer with a 13th Gen Intel(R) Core(TM) i9-13900H CPU, and $32$ GB of RAM.
Figure 4: Comparison of median execution times for DeVAIC, Bandit, CodeQL, Semgrep, and PyT for the analysis of all the $500$ codes generated by the $4$ models.
We compared the execution time taken by DeVAIC with the other solutions employed. Figure 4 shows the median execution time of each tool in analyzing $125$ code snippets from the four models. Bandit is the faster ($0.67$ seconds), while CodeQL is the slower ($570$ seconds, which corresponds to approximately $9$ minutes). PyT and Semgrep employed $253.40$ and $123.67$ seconds, respectively ($2$ and $4$ minutes approximately). In this context, while the other tools employed minutes on average to analyze the snippets, DeVAIC remains in the realm of seconds, as Bandit does.
Furthermore, we deeply inspected the distribution of the times required by DeVAIC to execute the detection rules on every code snippet. Despite the different lengths of the analyzed code snippets, Figure 5 shows that the tool exhibited an average execution time of $0.16$ seconds (with a median of $0.14$, a maximum of $0.59$, and a minimum of $0.10$). The figure highlights that for most of the snippets ($169$ in total, $\sim 34\%$), the execution time falls within the range of $0.10$ to $0.13$ seconds.
Figure 5: Occurrences of execution times taken by DeVAIC for scanning and evaluating individual code snippets generated by the 4 models.
Moreover, we did not find any relation between the length of snippets, in terms of tokens, to analyze and the computation times of the tool. Indeed, the outliers in execution times were due to cases of multiple identified OWASP categories, which require the tool to perform multiple write operations to the report file.
Figure 6: Execution times of DeVAIC plotted against the cumulative number of tokens in code snippets. Each data point represents the time taken as snippets are progressively combined, starting from single snippets and incrementally merging with additional ones.
Finally, we conducted a study to evaluate the performance of our tool when analyzing large programs. To increase the code complexity, we concatenated the snippets generated by Gemini (we could have chosen the code generated by any other AI model as well) to create a new set of code characterized by $125$ snippets with an incremental number of tokens, starting from $73$ for the first snippet to $6892$ for the last one.
Figure 6 shows the DeVAIC’s performance in evaluating the new code collection. Analyzing snippets with more tokens led to longer execution times. While the initial snippet, comprising $73$ tokens, only required $0.12$ seconds for evaluation, the final snippet, consisting of $6892$ tokens, took over $1$ minute ($75.23$ seconds) to process. Furthermore, we applied a second-degree polynomial curve to fit the data, obtaining an $R^{2}$ value of $0.984$. This value suggests a strong fit between the predicted execution times and the actual data and implies that the relationship between the snippet length and execution time is not linear but rather quadratic, with execution times increasing at a rate that accelerates with the number of tokens.
The polynomial relationship might be due to various factors inherent in code analysis processes, such as increased memory allocation, more intensive parsing, and a higher number of computational operations needed to process and analyze larger codebases.
6 Threats to Validity
Rule Creation: The creation of detection rules based on regular expressions could introduce bias, as these rules are derived from patterns observed within a limited set of vulnerable code snippets. This limitation may result in rules that are either too specific or too general, affecting the tool’s accuracy in real-world scenarios. To mitigate this threat, we employed a diverse dataset of vulnerabilities covering a broad range of CWEs and OWASP categories. We also conducted iterative refinements of our rules by testing them on separate validation sets to ensure they accurately capture the intended patterns without being overly broad or narrow.
AI Models: The selection of AI models for evaluating DeVAIC may introduce bias, as the performance of these models can vary significantly. This variation could inadvertently affect the perceived effectiveness of DeVAIC. However, we remark that we selected four widely used AI code generation models, representing a range of underlying technologies and training datasets. This diversity helps ensure that our evaluation encompasses a variety of code-generation behaviours and vulnerabilities.
Metrics: The use of Precision, Recall, $F_{1}$ Score, and Accuracy as metrics relies on correctly classifying vulnerabilities. Any misclassification could impact these metrics and the interpretation of DeVAIC effectiveness. To enhance the reliability of our classification, we employed a multi-researcher approach, where multiple experts independently assessed the vulnerabilities before reaching a consensus. This process reduces the risk of subjective bias and ensures a more accurate classification of true and false positives/negatives.
Generalizability: The effectiveness of DeVAIC is currently limited to Python code. This focus raises questions about the tool’s applicability to other programming languages that may have different syntax, semantics, and common vulnerabilities. While DeVAIC is initially designed for Python, the methodology used to create detection rules based on regular expressions and patterns observed in vulnerable code can be adapted to other languages. We are currently investigating this methodology for the extraction of vulnerable implementation patterns and rule creation in the C/C++ language. Adapting DeVAIC to C/C++ requires additional considerations due to the complexity and variety in how equivalent functionalities are implemented in these languages compared to Python (e.g., what can be accomplished with a single Python instruction might require several more complex instructions in C/C++). This complexity increases the difficulty of pattern identification and necessitates the development of more sophisticated detection rules. Future work will involve extending DeVAIC’s capabilities to additional languages and leveraging domain experts to ensure the accuracy and relevance of new rules.
7 Conclusion
In this work, we introduced DeVAIC, a tool that implements a set of detection rules to identify vulnerabilities in the Python code. The tool is designed to overcome the limitation of static analysis tools since it does not require complete programs, making it particularly suitable for AI-generated code. To define rules, we extracted vulnerable code from public datasets and, after grouping the code according to their OWASP categories and similarity, we found common patterns by standardizing the code snippets and using the LCS. We evaluated DeVAIC on code generated by four public AI-code generators. The results outline the tool’s ability to detect vulnerabilities, showing an average $F_{1}$ Score and Accuracy both at $94\%$, overcoming the performance of other state-of-the-art solutions used as a baseline for the evaluation, with a limited computational cost.
Future work aims to extend the number of program languages to analyze, enhancing the set of detection rules available. Moreover, we also aim to enrich the list of CWEs covered by the tool.